# Reciprocity formulas for certain generalized Hardy sums

### 05 Jan 2024 v3

@@ Abstract.
In this paper, we establish some reciprocity formulas for certain generalized Hardy sums by using the Fourier series technique and some properties of the periodic zeta function and Lerch zeta function. It turns out that one of Hardy’s reciprocity theorems is deduced as a special case.
@@
Keywords. quasi-periodic Euler functions; Bernoulli functions; Hardy sums; Hurwitz and Lerch zeta functions; Reciprocity formulas
@@ 1. Introduction
Let $\mathbb{N}$ be the set of positive integers, $\mathbb{N}_{0}$ the set of non-negative integers, $\mathbb{Z}$ the set of integers, $\mathbb{R}$ the set of real numbers, and $\mathbb{C}$ the set of complex numbers. Denote by $\{x\}$ the fractional part of $x\in\mathbb{R}$, and $((x))$ the sawtooth function given by
@@
where $s_{5}(a,b)$ is the Hardy sums given for $a,b\in\mathbb{Z}$ with $b\not=0$ by
@@
in which $[x]$ is the floor function satisfying $[x]=x-\{x\}$ for $x\in\mathbb{R}$. Just as Dedekind [12] deduced the reciprocity formula (1.2) from his transformation formula of the logarithm of the Dedekind eta function, Berndt [5] in 1978 used his transformation formulas of the logarithms of the classical theta-functions to prove three of Hardy’s reciprocity theorems. Goldberg [13] in 1981 used Berndt’s [5] transformation formulas to prove the reciprocity formula (1.3). In fact, Goldberg [13] also presented some three-term and mixed three-term reciprocity formulas for those sums considered by Hardy [14] and Berndt [5]. For some different proofs and extensions of Hardy’s reciprocity theorems, one is referred to [6, 7, 8, 11, 23, 27]. We also note that another interesting arithmetic properties for different types of Hardy sums have been studied in [19, 20, 22, 26, 29, 30].
The object of the present paper is to give some new extensions of the reciprocity formula (1.3). To do so, we consider the following generalized Hardy sums
@@
in which $E_{n}(x)$ is the Euler polynomial of degree $n$, $B_{n}(x)$ is the Bernoulli polynomial of degree $n$. By making use of the Fourier series technique and some properties of the periodic zeta function and Lerch zeta function, we establish some reciprocity formulas for the generalized Hardy sums (1.5). As immediate consequences of the results presented here, we rediscover the reciprocity formula (1.3).
@@
where $\mathrm{sgn}(x)$ indicates the sign of $x$ given for $x\in\mathbb{R}$ with $x\not=0$ by $\mathrm{sgn}(x)=x/|x|$, $\delta_{l,k}$ is the Kronecker delta function given for $l,k\in\mathbb{N}_{0}$ by $\delta_{l,k}=1$ or 00 according to $l=k$ or $l\not=k$, $\delta_{\mathbb{Z}}(x)=1$ or 00 according to $x\in\mathbb{Z}$ or $x\not\in\mathbb{Z}$, $(a,b)$ denotes the greatest common divisor of two integers $a$ and $b$.
@@ 2. Some auxiliary results
This ensures that the $n$-th quasi-periodic Euler function $\overline{E}_{n}(x)$ defined in (1.6) can be given for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [3, Theorem 1.2] or [21, Theorem 2.2])
@@
the $n$-th Bernoulli function $\overline{B}_{n}(x)$ defined in (1.7) can be given for $n\in\mathbb{N}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [2, Theorem 12.19] or [3, Theorem 1.1] or [21, Theorem 2.1])
@@
$a^{n-1}\sum_{r=0}^{a-1}\overline{B}_{n}\biggl{(}x+\frac{r}{a}\biggl{)}=\overline{B}_{n}(ax),$ (2.4)
@@
Since $\cot(a)$ has the following expression in partial fractions (see, e.g., [1, p. 75] or [25, p. 327])
@@
We now give a more general result of Lemma 2.5 as follows.
@@
The proof of Theorem 1.3. It is easily seen from Lemma 2.2 and (3.14) that for $j\in\mathbb{N}$, $c\in\mathbb{Z}$, $x\in\mathbb{R}$ with $c\not=0$,

### 15 Jan 2024 v5

# Reciprocity formulas for certain generalized Hardy-Berndt sums

@@ Abstract.
In this paper, we establish some reciprocity formulas for certain generalized Hardy-Berndt sums by using the Fourier series technique and some properties of the periodic zeta function and the Lerch zeta function. It turns out that one of Hardy’s reciprocity theorems is deduced as a special case.
@@
Keywords. quasi-periodic Euler functions; Bernoulli functions; Hardy-Berndt sums; Hurwitz and Lerch zeta functions; Reciprocity formulas
@@ 1. Introduction
Let $\mathbb{N}$ be the set of positive integers, $\mathbb{N}_{0}$ the set of non-negative integers, $\mathbb{Z}$ the set of integers, $\mathbb{R}$ the set of real numbers, and $\mathbb{C}$ the set of complex numbers. Denote by $\{x\}$ the fractional part of $x\in\mathbb{R}$, and $((x))$ the sawtooth function (also called the first Bernoulli function) given by
@@
where $s_{5}(a,b)$ is the Hardy sums (also known the Hardy-Berndt sums) given for $a,b\in\mathbb{Z}$ with $b\not=0$ by
@@
in which $[x]$ is the floor function satisfying $[x]=x-\{x\}$ for $x\in\mathbb{R}$. Just as Dedekind [12] deduced the reciprocity formula (1.2) from his transformation formula of the logarithm of the Dedekind eta function, Berndt [5] in 1978 used his transformation formulas of the logarithms of the classical theta-functions to prove three of Hardy’s reciprocity theorems. Goldberg [13] in 1981 used Berndt’s [5] transformation formulas to prove the reciprocity formula (1.3). In fact, Goldberg [13] also presented some three-term and mixed three-term reciprocity formulas for those sums considered by Hardy [14] and Berndt [5]. For some different proofs and extensions of Hardy’s reciprocity theorems, one is referred to [6, 7, 8, 11, 20, 23].
The object of the present paper is to give some new extensions of the reciprocity formula (1.3). To do so, we consider the following generalized Hardy-Berndt sums
@@
in which $E_{n}(x)$ is the Euler polynomial of degree $n$, $B_{n}(x)$ is the Bernoulli polynomial of degree $n$. By using the Fourier series technique and some properties of the periodic zeta function and the Lerch zeta function, we establish some reciprocity formulas for the generalized Hardy-Berndt sums (1.5). As immediate consequences of the results presented here, we rediscover the reciprocity formula (1.3).
@@
where $\mathrm{sgn}(x)$ indicates the sign of $x$ given for $x\in\mathbb{R}\setminus\{0\}$ by $\mathrm{sgn}(x)=x/|x|$, $\delta_{l,k}$ is the Kronecker delta function given for $l,k\in\mathbb{N}_{0}$ by $\delta_{l,k}=1$ or 00 according to $l=k$ or $l\not=k$, $\delta_{\mathbb{Z}}(x)=1$ or 00 according to $x\in\mathbb{Z}$ or $x\not\in\mathbb{Z}$, $(a,b)$ denotes the greatest common divisor of two integers $a,b$.
@@
It should be noted that replacing $a$ by $c$ and $c$ by $a$ in Corollary 1.4, and then using the relations shown in [20, Equations (3.7), (3.8) and (3.9)], we can reobtain Goldberg’s [13, Theorem 5.6] mixed three-term reciprocity formula.
@@ 2. Some auxiliary results
This ensures that the $n$-th quasi-periodic Euler function $\overline{E}_{n}(x)$ defined in (1.6) can be given for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [3, Theorem 1.2] or [19, Theorem 2.2])
@@
the $n$-th Bernoulli function $\overline{B}_{n}(x)$ defined in (1.7) can be given for $n\in\mathbb{N}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [2, Theorem 12.19] or [3, Theorem 1.1] or [19, Theorem 2.1])
@@
$a^{n-1}\sum_{r=0}^{a-1}\overline{B}_{n}\biggl{(}x+\frac{r}{a}\biggl{)}=\overline{B}_{n}(ax).$ (2.4)
@@
Since $\cot(a)$ has the following expression in partial fractions (see, e.g., [1, p. 75] or [22, p. 327])
@@
The proof of Theorem 1.3. It is easily seen from Lemma 2.2 and (3.14) that for $j\in\mathbb{N}$, $c\in\mathbb{Z}\setminus\{0\}$, $x\in\mathbb{R}$,